5 minute scans were taken on 18nov09 and 19nov09 to
check alfa's stability. the setup was:
- az=285,za=11, alfa_angle=0 18nov09, angle=19 19nov09
- .1 sec integrations for 300 seconds.
- Mock spectrometer with two 172 MHz bands of 8192 channels each
centered at 1300 and 1450 MHz.
- a/d blanking enabled.
The data processing was:
- correct for a/d blanking
- exclude channels with rfi by computing rms/channel for 5 minutes,
linear fit, and throw out outliers.
- Compute power over the good channels.
The plots show the total power vs time
for the two days (.ps)
- page 1: 18nov09
- page2: 19nov09
- frames 1-2 polA 1300 and 1450 MHz bands
- frames 3-4 polB 1300 and 1450 MHz bands.
- Colors: bottom to top is beam 0 to 6 (black=bm0,red=bm1 etc..)
- bm2a 1450 is noisy on 18nov09 but not on 19nov09 so who knows..
- bm5a: both 1300,1450 bands and both day. It is noisier than
the other beams.
- bm1a: 1450 band only, both days wonders around.
- This was also seen on the led recovery test.